3) Prove by induction that for all natural numbers n:


(c) 1/(1)(2) + 1/(2)(3) + · · · +1/(n)(n + 1) = n/(n + 1)



1
Expert's answer
2022-01-10T15:13:25-0500

Let "P(n)" be the statement that "\frac{1}{1\cdot 2}+\frac{1}{2\cdot 3}+…+\frac{1}{n\cdot(n+1)}=\frac{n}{n+1}" .


1) For "n=1" we have that "\frac{1}{1\cdot 2}=\frac{1}{2}=\frac{1}{1+1}" . So, "P(1)" is true.


2) Suppose that "P(k)" holds, that is, "\frac{1}{1\cdot 2}+\frac{1}{2\cdot 3}+…+\frac{1}{k\cdot(k+1)}=\frac{k}{k+1}" .


"\frac{1}{1\cdot 2}+\frac{1}{2\cdot 3}+…+\frac{1}{k\cdot(k+1)}+\frac{1}{(k+1)(k+2)}=\frac{k}{k+1}+\frac{1}{(k+1)(k+2)}=\frac{k(k+2)+1}{(k+1)(k+2)}=\frac{k^2+2k+1}{(k+1)(k+2)}=\frac{(k+1)^2}{(k+1)(k+2)}=\frac{k+1}{k+2}"


We have that "\frac{1}{1\cdot 2}+\frac{1}{2\cdot 3}+…+\frac{1}{(k+1)\cdot(k+2)}=\frac{k+1}{(k+1)+1}" .


Therefore, "P(k+1)" holds.


By the Principle of Mathematical Induction, for all "n\in\mathbb{N}:\" "\frac{1}{1\cdot 2}+\frac{1}{2\cdot 3}+…+\frac{1}{n\cdot(n+1)}=\frac{n}{n+1}"